OpenStudy (anonymous):

What is a perfect square and the difference between to squares. (polynomials & algebra 1)

OpenStudy (anonymous):

an integer n is called a perfect square if n = k^2 for some integer k

OpenStudy (anonymous):

I am sorry could you explain that a little more.

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

is 4 a perfect square?

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I am going to take my module 7 DBA in algebra one this afternoon. It says Perfect square trinomial... (a+b)^2=a^2+2ab+b^2.

OpenStudy (anonymous):

sorry i never took algebra 1 :/

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@NemesisFeeDz

OpenStudy (anonymous):

@surjithayer you still there???

OpenStudy (triciaal):

difference between 2 squares difference what you get when you subtract a^2 - b^2 trinomial 3 terms a^2 - b^2 = (a-b)(a+b) = a^2 -ab + ab + b^2 but the -ab and the + ab cancel out (a-b)^2 = (a-b)(a-b)= a^2 -2ab -b^2

OpenStudy (anonymous):

Oh okay thank you :]

OpenStudy (triciaal):

a perfect square is what you get when you multiply the entity by itself like previous response above 4 is a perfect square because the square root is an integer can be expressed as a ratio

OpenStudy (anonymous):

A perfect square trinomial.

OpenStudy (anonymous):

?

OpenStudy (triciaal):